OpenStudy (anonymous):

Write the inverse variation equation of the following. Y varies inversely with x and y =10 when x = 50. HELP?

OpenStudy (anonymous):

Subtract 10 from y to set y=0. Since x and t are inversely variable, add 10 to x to make x=60 So y=60-x Or x=60-y, whichever you prefer.

OpenStudy (anonymous):

Thank you! Glad I could help! :-)